OpenStudy (anonymous):

Calculate ∆H for the reaction of nitrous oxide with oxygen. 2NO(g) + O2(g) --> 2NO2(g) ∆Hf0 NO(g) =90.37 kJ/mol; ∆Hf0 NO2 (g) =33.85 kJ/mol)

OpenStudy (abmon98):

∆H products-∆H reactants= (33.85*2)KJ-(90.37*2)KJ=67.7-180.74=-113.04 reaction is exothermic